# SoundnessBench: A Soundness Benchmark for Neural Network Verifiers

### v2

@@ 3.2 Overall Pipeline
Our benchmark is constructed with a number of different settings. Each setting is identified by: 1) A dataset that provides original examples ${\bm{x}}_{0}$ and ground-truth labels $y$; 2) A neural network architecture for model $f$; 3) The perturbation size $\epsilon$}. For each setting, our pipeline proceeds as follows: For each example from the original dataset, we generate a random target class. We aim to make half of these examples unverifiable with hidden counterexamples (i.e., unverifiable instances), by generating a random pre-defined counterexample. The other half are regular examples without pre-defined counterexamples, so that these examples are potentially verifiable and NN verifier developers do not simply know all the examples are unverifiable. We then train a model with both an adversarial training objective and an additional objective for producing counterexamples. After the training, for the examples with pre-defined counterexamples, we filter out those with counterexamples that can be found by a stronger adversarial attack, or where the pre-defined counterexamples do not end up becoming counterexamples. We discuss these stages in more detail below. We now focus on a given setting with model $f$ and perturbation size $\epsilon$}, and we omit $f$ and $\epsilon$} when describing instances below. Figure˜1 provides an overview.
@@ 3.3 Data generation
In these instances, ${\bm{x}}_{0}$ and $y$ come from an original dataset, where the first $n$ examples are used to construct ${\mathbb{D}}_{\text{unv}}$ while the remaining $n$ examples are used for ${\mathbb{D}}_{\text{regular}}$. For each unverifiable instance $({\bm{x}}^{(\text{unv},i)}_{0},y^{(\text{unv},i)})\in{\mathbb{D}}_{\text{unv}}~(1\leq i\leq n)$, we generate a pre-defined counterexample $({\bm{x}}_{\text{CEX}}^{(i)},y_{\text{CEX}}^{(i)})$, where ${\bm{x}}_{\text{CEX}}^{(i)}={\bm{x}}^{(\text{unv},i)}_{0}+\delta_{\text{CEX}}^{(i)}$, such that $\delta_{\text{CEX}}^{(i)}$ is the perturbation and $y_{\text{CEX}}^{(i)}$ is the target label. $\delta_{\text{CEX}}^{(i)}$ is randomly generated from $([-\epsilon,-r\cdot{\epsilon}]\bigcup[r\cdot{\epsilon},{\epsilon}])^{d}~(0\leq r<1)$ satisfying $r\cdot{\epsilon}\leq\|\delta_{\text{CEX}}^{(i)}\|_{\infty}\leq{\epsilon}$, and $y_{\text{CEX}}^{(i)}$ is randomly generated from $\{j\mid 1\leq j\leq K,j\neq y^{(\text{unv},i)}\}$ satisfying $y_{\text{CEX}}^{(i)}\neq y^{(\text{unv},i)}$. Since we aim to ensure ${\bm{x}}^{(\text{unv},i)}_{0}$ and ${\bm{x}}^{(i)}_{\text{CEX}}$ have different labels, we make them separated by setting $\|\delta_{\text{CEX}}^{(i)}\|_{\infty}\geq r\cdot{\epsilon}$ with a hyperparameter $r~(0\leq r<1)$.
@@ 3.4 Training
In the first objective regarding Eq. (4), for each unverifiable instance $({\bm{x}}^{(\text{unv},i)}_{0},y^{(\text{unv},i)})~(1\leq i\leq n)$ with the pre-defined counterexample $({\bm{x}}_{\text{CEX}}^{(i)},y_{\text{CEX}}^{(i)})$, we aim to make the class predicted from output logits $f({\bm{x}}_{\text{CEX}}^{(i)})$ match target label $y_{\text{CEX}}^{(i)}$, i.e., $\operatorname*{arg\,max}_{1\leq j\leq K}f_{j}({\bm{x}}_{\text{CEX}}^{(i)})=y_{\text{CEX}}^{(i)}.$ This can be achieved by a standard cross-entropy (CE) loss. However, it tends to maximize $f_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})$ and minimize $f_{y^{(\text{unv},i)}}({\bm{x}}^{(\text{unv},i)})$ without a cap on the margin between them. We find that making $f_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})$ substantially greater than $f_{y^{(\text{unv},i)}}({\bm{x}}^{(\text{unv},i)})$ tends to make the counterexamples more obvious to adversarial attacks and thus makes our training with the second objective difficult. Therefore, we design a margin objective instead to achieve $0\leq f_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})-f_{y^{(\text{unv},i)}}({\bm{x}}^{(\text{unv},i)})\leq\lambda$ with a maximum desired margin $\lambda$:
${\mathcal{L}}_{\text{CEX}}=\frac{1}{n}\sum_{i=1}^{n}\max\{0,f_{y^{(\text{unv},i)}}({\bm{x}}^{(\text{unv},i)})-f_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})+\lambda)\}.$ (9)
@@
Our second objective is adversarial training as defined in Eq. (2), applied to the entire dataset $\mathbb{D}$} with both unverifiable instances and regular instances. However, since by default a new perturbation $\delta^{(i)}~(1\leq i\leq 2n)$ is generated at each training epoch for every example, we find that the training struggles to converge, if we only use the new perturbation. This is because training on a new perturbation can often break the model on other perturbations including those used in previous epochs, when we have a conflicting objective for pre-defined counterexamples. To mitigate this issue, we propose a perturbation sliding window, where we use multiple perturbations generated from the most recent epochs. For each instance $({\bm{x}}_{0}^{(i)},y^{(i)})$, we maintain a window ${\mathbb{W}}^{(i)}$ of perturbations generated in the most recent $w$ epochs: ${\mathbb{W}}^{(i)}=\{\delta_{1}^{(i)},\delta_{2}^{(i)},\cdots,\delta_{w}^{(i)}\}$, where $w$ is the maximum length of the sliding window (there can be fewer than $w$ elements in ${\mathbb{W}}^{(i)}$ in the first w1w-1 epochs). Using this window, our objective that utilizes all the perturbations in the sliding window is:
${\mathcal{L}}_{\text{adv}}={\mathcal{L}}_{\text{adv, window}}=\frac{1}{2n}\sum_{1\leq i\leq 2n}\frac{1}{|{\mathbb{W}}^{(i)}|}\sum_{\delta\in{\mathbb{W}}^{(i)}}{\mathcal{L}}_{\text{CE}}(f({\bm{x}}_{0}^{(i)}+\delta),y^{(i)}).$ (10)
@@
Suppose model $f$ is parameterized by parameters $\theta$, denoted as $f_{\theta}$, and our two loss functions both depend on $\theta$, denoted as ${\mathcal{L}}_{\text{CEX}}(\theta)$ and ${\mathcal{L}}_{\text{adv}}(\theta)$. Then our overall loss function combines the two objectives as:
${\mathcal{L}}(\theta)={\mathcal{L}}_{\text{CEX}}(\theta)+{\mathcal{L}}_{\text{adv}}(\theta),$ (11)
and we optimize model parameters by solving $\operatorname*{arg\,min}_{\theta}{\mathcal{L}}(\theta)$ using a gradient-based optimizer.
@@ 3.5 Benchmark Design
Name Model Architecture Activation Function CNN 1 Conv Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU CNN 2 Conv Conv $5\times 3\times 3$, Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU CNN 3 Conv Conv $5\times 3\times 3$, Conv $10\times 3\times 3$, Conv 20 $3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU CNN AvgPool Conv $10\times 3\times 3$, AvgPool $3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU MLP 4 Hidden FC 100, FC 1000, FC 1000, FC 1000, FC 20, FC 2 ReLU MLP 5 Hidden FC 100, FC 1000, FC 1000, FC 1000, FC 1000, FC 20, FC 2 ReLU CNN Tanh Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 Tanh CNN Sigmoid Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 Sigmoid ViT Modified ViT (Shi et al., 2025) with patch size $1\times 1$, 2 attention heads and embedding size 16 ReLU
@@
We focus on using relatively small models and low-dimensional synthetic data. Since the goal of our benchmark is to test the soundness of NN verifiers rather than benchmark performance, our models and data do not have to be large-scale or as practical as real-world ones. We may draw an analogy to fuzzing techniques in software engineering, where simple and synthetic but specifically constructed inputs, rather than complicated real-world inputs, are used for testing software and identifying bugs. Additionally, large-scale models and data can significantly increase the difficulty for NN verifiers, and NN verifiers can fail to provide meaningful verification results or simply fail to support complicated cases, without exposing bugs. This situation is more pronounced when models are trained by our method with pre-defined counterexamples (models and data at our small scale are already difficult for current NN verifiers, as shown in Table˜6). Therefore, by using relatively small models and data, we allow more verifiers to be tested against our benchmark to maximize bug exposure.
@@
Our dataset of original examples $({\bm{x}}_{0}^{(1)},y^{(1)}),({\bm{x}}_{0}^{(2)},y^{(2)}),\cdots,({\bm{x}}_{0}^{(n)},y^{(n)})$ is synthetic and randomly generated following uniform distributions, where ${\bm{x}}_{0}^{(i)}$ is sampled from $1,1]^{d}$ and $y^{(i)}$ is sampled from $\{0,1\}$ for $1\leq i\leq 2n$. We use various input sizes and perturbation radii, as shown in LABEL:tab:training_results. Specifically, for MLP models, we use 1D data with input size $d=10$; for other models, we use 2D data with input sizes $d=1\times 5\times 5$ and $d=3\times 5\times 5$, respectively, where $1$ and $3$ are the numbers of input channels considered by CNNs and ViTs. For the perturbation radii, we use ${\epsilon}=0.2$ and ${\epsilon}=0.5$. For each model with a particular perturbation radius and input size, we set $n=10$ as the number of instances and generate $10$ unverifiable instances $({\bm{x}}_{0}^{(\text{unv},i)},y^{(\text{unv},i)})\in{\mathbb{D}}_{\text{unv}}~(1\leq i\leq n)$, each with a corresponding pre-defined counterexample $({\bm{x}}_{\text{CEX}}^{(i)},y_{\text{CEX}}^{(i)})$, and $10$ regular instances $({\bm{x}}_{0}^{(\text{regular},i)},y^{(\text{regular},i)})\in{\mathbb{D}}_{\text{regular}}~(1\leq i\leq n)$. We set $r=0.98$ for $\|\delta_{\text{CEX}}^{(i)}\|_{\infty}\geq r\cdot{\epsilon}$, introduced in Section˜3.3, so that each pre-defined counterexample is close to the boundary of the input perturbation ball. During the data generation, to ensure that different examples are sufficiently separated, we require that there is no intersection between the perturbation balls of different examples and retry the generation until no intersection exists. After the training, for the unverifiable instances, we only keep those with hidden counterexamples that are not found by strong adversarial attacks conducted in Section˜4.2. Thus, there can be fewer than 10 unverifiable instances if any pre-defined counterexample is found by an adversarial attack.
@@ 4 Experiments
In this section, we conduct comprehensive experiments. We first build our SoundnessBench by our proposed method with implementation details provided in Section˜4.1, we evaluate the effectiveness of our training in Section˜4.2. We then use our benchmark to test the soundness of several state-of-the-art NN verifiers in Section˜4.3. Additionally, in Section˜4.4, we conduct an ablation study to justify the design choices for our training. We also provide additional results in the appendix about testing NN verifiers with synthetic bugs (Appendix˜A) and repeating our experiments with multiple random seeds (Appendix˜B).

### v3

@@ 3.2 Overall Pipeline
Our benchmark is constructed with a number of different settings. Each setting is identified by: 1) A dataset that provides original examples ${\bm{x}}_{0}$ and ground-truth labels $y$; 2) A neural network architecture for model $\bm{f}$}; 3) The perturbation size $\epsilon$}. For each setting, our pipeline proceeds as follows: For each example from the original dataset, we generate a random target class. We aim to make half of these examples unverifiable with hidden counterexamples (i.e., unverifiable instances), by generating a random pre-defined counterexample. The other half are regular examples without pre-defined counterexamples, so that these examples are potentially verifiable and NN verifier developers do not simply know all the examples are unverifiable. We then train a model with both an adversarial training objective and an additional objective for producing counterexamples. After the training, for the examples with pre-defined counterexamples, we filter out those with counterexamples that can be found by a stronger adversarial attack, or where the pre-defined counterexamples do not end up becoming counterexamples. We discuss these stages in more detail below. We now focus on a given setting with model $\bm{f}$} and perturbation size $\epsilon$}, and we omit $\bm{f}$} and $\epsilon$} when describing instances below. Figure˜1 provides an overview.
@@ 3.3 Data generation
In these instances, ${\bm{x}}_{0}$ and $y$ come from an original dataset, where the first $n$ examples are used to construct ${\mathbb{D}}_{\text{unv}}$ while the remaining $n$ examples are used for ${\mathbb{D}}_{\text{regular}}$. For each unverifiable instance $({\bm{x}}^{(\text{unv},i)}_{0},y^{(\text{unv},i)})\in{\mathbb{D}}_{\text{unv}}~(1\leq i\leq n)$, we generate a pre-defined counterexample $({\bm{x}}_{\text{CEX}}^{(i)},y_{\text{CEX}}^{(i)})$, where ${\bm{x}}_{\text{CEX}}^{(i)}={\bm{x}}^{(\text{unv},i)}_{0}+{\bm{\delta}}_{\text{CEX}}^{(i)}$, such that ${\bm{\delta}}_{\text{CEX}}^{(i)}$ is the perturbation and $y_{\text{CEX}}^{(i)}$ is the target label. ${\bm{\delta}}_{\text{CEX}}^{(i)}$ is randomly generated from $([-\epsilon,-r\cdot{\epsilon}]\bigcup[r\cdot{\epsilon},{\epsilon}])^{d}~(0\leq r<1)$ satisfying $r\cdot{\epsilon}\leq\|{\bm{\delta}}_{\text{CEX}}^{(i)}\|_{\infty}\leq{\epsilon}$, and $y_{\text{CEX}}^{(i)}$ is randomly generated from $\{j\mid 1\leq j\leq K,j\neq y^{(\text{unv},i)}\}$ satisfying $y_{\text{CEX}}^{(i)}\neq y^{(\text{unv},i)}$. Here, since we aim to ensure ${\bm{x}}^{(\text{unv},i)}_{0}$ and ${\bm{x}}^{(i)}_{\text{CEX}}$ have different labels, we make them separated by setting $\|{\bm{\delta}}_{\text{CEX}}^{(i)}\|_{\infty}\geq r\cdot{\epsilon}$ with a hyperparameter $r~(0\leq r<1)$ for the minimum distance, where $r$ is set to a value close to 1. This is similar to how zimmermann2022increasing controlled the distance and generated perturbed examples near the boundary of the ${\mathcal{B}}({\bm{x}}_{0},{\epsilon})$ perturbation ball when creating their dataset for planting counterexamples.
@@ 3.4 Training
In the first objective regarding Eq. (4), for each unverifiable instance $({\bm{x}}^{(\text{unv},i)}_{0},y^{(\text{unv},i)})~(1\leq i\leq n)$ with the pre-defined counterexample $({\bm{x}}_{\text{CEX}}^{(i)},y_{\text{CEX}}^{(i)})$, we aim to make the class predicted from output logits ${\bm{f}}({\bm{x}}_{\text{CEX}}^{(i)})$ match target label $y_{\text{CEX}}^{(i)}$, i.e., $\operatorname*{arg\,max}_{1\leq j\leq K}{\bm{f}}_{j}({\bm{x}}_{\text{CEX}}^{(i)})=y_{\text{CEX}}^{(i)}.$ This can be achieved by a standard cross-entropy (CE) loss. However, it tends to maximize ${\bm{f}}_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})$ and minimize ${\bm{f}}_{y^{(\text{unv},i)}}({\bm{x}}_{\text{CEX}}^{(i)})$ without a cap on the margin between them. We find that making ${\bm{f}}_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})$ substantially greater than ${\bm{f}}_{y^{(\text{unv},i)}}({\bm{x}}_{\text{CEX}}^{(i)})$ tends to make the counterexamples more obvious to adversarial attacks and thus makes our training with the second objective difficult. Therefore, we design a margin objective instead to achieve $0\leq{\bm{f}}_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})-{\bm{f}}_{y^{(\text{unv},i)}}({\bm{x}}_{\text{CEX}}^{(i)})\leq\lambda$ with a maximum desired margin $\lambda$:
${\mathcal{L}}_{\text{CEX}}=\frac{1}{n}\sum_{i=1}^{n}\max\{0,{\bm{f}}_{y^{(\text{unv},i)}}({\bm{x}}_{\text{CEX}}^{(i)})-{\bm{f}}_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})+\lambda)\}.$ (10)
This is different from zimmermann2022increasing as they only trained a linear classification head to separate the frozen features of perturbed examples inside the perturbation ball and on the boundary of the perturbation ball, thereby making the examples on the boundary become counterexamples. In contrast, we train the whole model, as we additionally have a more challenging objective of making the counterexamples hidden against adversarial attacks (achieved via the adversarial training objective described below), and we achieve the objective for the pre-defined counterexamples here with a margin objective.
@@
Our second objective is adversarial training as defined in Eq. (2), applied to the entire dataset $\mathbb{D}$} with both unverifiable instances and regular instances. However, since by default a new perturbation ${\bm{\delta}}^{(i)}~(1\leq i\leq 2n)$ is generated at each training epoch for every example, we find that the training struggles to converge, if we only use the new perturbation. This is because training on a new perturbation can often break the model on other perturbations including those used in previous epochs, when we have a conflicting objective for pre-defined counterexamples. To mitigate this issue, we propose a perturbation sliding window, where we use multiple perturbations generated from the most recent epochs. For each instance $({\bm{x}}_{0}^{(i)},y^{(i)})$, we maintain a window ${\mathbb{W}}^{(i)}$ of perturbations generated in the most recent $w$ epochs: ${\mathbb{W}}^{(i)}=\{{\bm{\delta}}_{1}^{(i)},{\bm{\delta}}_{2}^{(i)},\cdots,{\bm{\delta}}_{w}^{(i)}\}$, where $w$ is the maximum length of the sliding window (there can be fewer than $w$ elements in ${\mathbb{W}}^{(i)}$ in the first w1w-1 epochs). Using this window, our objective that utilizes all the perturbations in the sliding window is:
${\mathcal{L}}_{\text{adv}}={\mathcal{L}}_{\text{adv, window}}=\frac{1}{2n}\sum_{1\leq i\leq 2n}\frac{1}{|{\mathbb{W}}^{(i)}|}\sum_{{\bm{\delta}}\in{\mathbb{W}}^{(i)}}{\mathcal{L}}_{\text{CE}}({\bm{f}}({\bm{x}}_{0}^{(i)}+{\bm{\delta}}),y^{(i)}).$ (11)
This differs significantly from zimmermann2022increasing which did not have adversarial training to make counterexamples hidden.
@@
Suppose model $\bm{f}$} is parameterized by parameters $\bm{\theta}$}, denoted as ${\bm{f}}_{\bm{\theta}}$, and our two loss functions both depend on $\bm{\theta}$}, denoted as ${\mathcal{L}}_{\text{CEX}}({\bm{\theta}})$ and ${\mathcal{L}}_{\text{adv}}({\bm{\theta}})$. Then our overall loss function combines the two objectives as:
${\mathcal{L}}({\bm{\theta}})={\mathcal{L}}_{\text{CEX}}({\bm{\theta}})+{\mathcal{L}}_{\text{adv}}({\bm{\theta}}),$ (12)
and we optimize model parameters by solving $\operatorname*{arg\,min}_{\bm{\theta}}{\mathcal{L}}({\bm{\theta}})$ using a gradient-based optimizer. The complete training algorithm is shown in Algorithm˜1 in Appendix A.
@@ 3.5 Benchmark Design
Name Model Architecture Activation Function CNN 1 Conv Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU CNN 2 Conv Conv $5\times 3\times 3$, Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU CNN 3 Conv Conv $5\times 3\times 3$, Conv $10\times 3\times 3$, Conv 20 $3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU CNN AvgPool Conv $10\times 3\times 3$, AvgPool $3\times 3$, FC 1000, FC 100, FC 20, FC 2 ReLU MLP 4 Hidden FC 100, FC 1000, FC 1000, FC 1000, FC 20, FC 2 ReLU MLP 5 Hidden FC 100, FC 1000, FC 1000, FC 1000, FC 1000, FC 20, FC 2 ReLU CNN Tanh Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 Tanh CNN Sigmoid Conv $10\times 3\times 3$, FC 1000, FC 100, FC 20, FC 2 Sigmoid ViT Modified ViT (shi2024genbab) with patch size $1\times 1$, 2 attention heads and embedding size 16 ReLU
@@
We focus on using relatively small models and low-dimensional synthetic data. Since the goal of our benchmark is to test the soundness of NN verifiers rather than benchmark performance, our models and data do not have to be large-scale or as practical as real-world ones. We may draw an analogy to fuzzing techniques in software engineering, where simple and synthetic but specifically constructed inputs, rather than complicated real-world inputs, are used for testing software and identifying bugs. Additionally, large-scale models and data can significantly increase the difficulty for NN verifiers, and NN verifiers can fail to provide meaningful verification results or simply fail to support complicated cases, without exposing bugs. Verifiers typically require relaxation, and on relatively larger settings, looser relaxation of other correct layers can compensate for the wrong bound computation, making the final verification more conservative. Although NN verifiers can often verify much larger models in other benchmarks, they depend on models being specifically trained, and our models trained with hidden counterexamples are more challenging to verify. As shown in Table˜6, models and data at our small scale are already difficult for current NN verifiers. Therefore, by using relatively small models and data, we allow more verifiers to be tested against our benchmark to maximize bug exposure. Nevertheless, we also show results on MNIST in Appendix E, where verification is much more difficult.
@@
Our dataset of original examples $({\bm{x}}_{0}^{(1)},y^{(1)}),({\bm{x}}_{0}^{(2)},y^{(2)}),\cdots,({\bm{x}}_{0}^{(n)},y^{(n)})$ is synthetic and randomly generated following uniform distributions, where ${\bm{x}}_{0}^{(i)}$ is sampled from $1,1]^{d}$ and $y^{(i)}$ is sampled from $\{0,1\}$ for $1\leq i\leq 2n$. We use various input sizes and perturbation radii, as shown in LABEL:tab:training_results. Specifically, for MLP models, we use 1D data with input size $d=10$; for other models, we use 2D data with input sizes $d=1\times 5\times 5$ and $d=3\times 5\times 5$, respectively, where $1$ and $3$ are the numbers of input channels considered by CNNs and ViTs. For the perturbation radii, we use ${\epsilon}=0.2$ and ${\epsilon}=0.5$. For each model with a particular perturbation radius and input size, we set $n=10$ as the number of instances and generate $10$ unverifiable instances $({\bm{x}}_{0}^{(\text{unv},i)},y^{(\text{unv},i)})\in{\mathbb{D}}_{\text{unv}}~(1\leq i\leq n)$, each with a corresponding pre-defined counterexample $({\bm{x}}_{\text{CEX}}^{(i)},y_{\text{CEX}}^{(i)})$, and $10$ regular instances $({\bm{x}}_{0}^{(\text{regular},i)},y^{(\text{regular},i)})\in{\mathbb{D}}_{\text{regular}}~(1\leq i\leq n)$. We set $r=0.98$ for $\|{\bm{\delta}}_{\text{CEX}}^{(i)}\|_{\infty}\geq r\cdot{\epsilon}$, introduced in Section˜3.3, so that each pre-defined counterexample is close to the boundary of the input perturbation ball. During the data generation, to ensure that different examples are sufficiently separated, we require that there is no intersection between the perturbation balls of different examples and retry the generation until no intersection exists. After the training, for the unverifiable instances, we only keep those with hidden counterexamples that are not found by strong adversarial attacks conducted in Section˜4.2. Thus, there can be fewer than 10 unverifiable instances if any pre-defined counterexample is found by an adversarial attack.
We compare settings with existing benchmarks in Appendix B.
@@ 4 Experiments
In this section, we conduct comprehensive experiments. We first build our SoundnessBench by our proposed method with implementation details provided in Section˜4.1, we evaluate the effectiveness of our training in Section˜4.2. We then use our benchmark to test the soundness of several state-of-the-art NN verifiers in Section˜4.3. Additionally, in Section˜4.4, we conduct an ablation study to justify the design choices for our training. We also provide additional results in the appendix about testing NN verifiers with synthetic bugs (Appendix C), repeating our experiments with multiple random seeds (Appendix D), and experimenting on MNIST (Appendix E).